Project

General

Profile

Actions

2017-08-16-LH2

Phone details

Present: KL, AN, PH, AK, PW, JT, CMacW, HN
Apologies: JB

Notes.

  1. Actions
    • AK/PW: test permissions of required connections HMI to PLC.
      • Stands: Will take a look at it today.
    • MT: to fit and test PV20 solenoid.
      • Stands: We didnt know if this had been done.
    • PW/MT: to test PV20 in sequence.
      • Stands: To be done once it is fittled.
    • JB: to prepare rota on 'limited call-out' basis.
      • Stands.
    • AK: to arrange meeting on archiver with JB.
      • Stands.
    • AN: Arrange final documentation review
      • Stands. Will be done on morning of the 31Aug17. Must incorporate some of points raised by J.Thomason (see below).
    • CMcW: Apply labels to pipes in loading bay area
      • Done.
    • CW: Secure crash barriers around charging enclosure to roadway.
      • Done.
    • AN: to apply ATEX zone notice to valve panel area.
      • Stands. Will be done when valve panel is complete.
    • AG: Check bonding of lightning conductor
      • Stands.
    • PW: to send PV conversion factors to AK. AK will send list of required PVs.
      • Done.
    • AK: to move monitoring GUI to ISIS control room. Audible alarm also required.
      • Stands. Remote session to HEPLNV154. Will implement today/tomorrow to make it live when documentation goes live.
    • PW: to write up required works for power supply for absorber heaters.
      • Stands. Will be implemented this week.
    • PW: to fit 4th 20mA supply to top level sensor.
      • Stands. Will be done this week.
    • CW: to fit Gates/chain to control access to valve panel Atex zone.
      • Done. Now need signage.
    • CMcW/HN: to modify Hall search to account for Atex zone barriers.
      • Stands.
    • CW: to find shifters for LH2 run 21st to 25th August.
      • Superceded.
    • PW: to add acknowledge-fault functionality to software.
      • Done.
  2. Monitoring and archival of data - PF/AK
    • Covered in previous meetings. Outstanding item is the discussion with JB and implementation of outcome of that meeting.
      • For alarm handled need to generate config file.
  3. Current technical
    • 1st elbow and T:
      • Done. P.Lester is welding final pieces. Will be fitted later today.
    • Merge:
      • Done.
    • Bellows bracing.
      • In hand: Rings have been fitted/welding. Stainless bars will arrive from RS today.
    • Measurements for remainder complete except final fit - Monday.
      • See above. Expect to be complete today. Tomorrow will bolt up and seal.
  4. Documentation
    • Met with J.Thomason yesterday the sstatus of the documentation and the approach to teh review. Will post on the wiki page and ask MICE/ISIS LH2 w/g to read. Final review will be on the morning pf the 31Aug17. Need ExVeritas verification of the LH2 pump enclosure on the MICE Hall roof. Agreed to brief ISIS Crew on crew instructions and instructions for Alarm Investigation Team.
    • Sign-off list is beginning to be populated.
    • Modified induction training to be implemented. SB must create necessary documentation.
      • Atex zone drawing - JT/JB: To issue document this week.
    • Items in H2 panel: all necessary traceablility has been estblished.
  5. Timetable
    • Expected outstanding on 23Aug17:
      • PV20 fitting and test;
  6. Run periods before 19th Sept.
    • Can cool-down as soon as system complete. Seek to fill with H2 as soon after the fial safety tour as possible.
      • AN: Will check that H2 gas is available. Will try and install in the bottle area.
      • LH2 sensor tests will be 11Sep17.
  7. Expert cover
    • Need to make rota.
  8. Next meeting
    • Wednesday 30Aug17; 11:00 MICE Meeting Room.
  9. AoB
    • HN: S.Boyd needs to make training happen and is recorded.

Summary of actions:

  • AK/PW: test permissions of required connections HMI to PLC.
  • MT: to fit and test PV20 solenoid.
  • PW/MT: to test PV20 in sequence.
  • JB: to prepare rota on 'limited call-out' basis.
  • AK: to arrange meeting on archiver with JB.
  • AN: Arrange final documentation review
  • AN: to apply ATEX zone notice to valve panel area.
  • AG: Check bonding of lightning conductor
  • AK: to move monitoring GUI to ISIS control room. Audible alarm also required.
  • PW: to write up required works for power supply for absorber heaters.
  • PW: to fit 4th 20mA supply to top level sensor.
  • CMcW/HN: to modify Hall search to account for Atex zone barriers.
  • SB: create documentation necessary to support shifter training.
  • JT/JB: To issue ATEX zoning document this week (i.e. with the other documentation).
  • AN: Will check that H2 gas is available to allow trial installation of gas pack in H2 locker area.

Files:

Updated by Long, Kenneth almost 7 years ago · 3 revisions